OpenStudy (anonymous):

write the percent as a fraction in simplest forn and as a decimal 15.5%

OpenStudy (agentjamesbond007):

percent is always over 100, so it would be 15.5/100 It might be in simplest form

OpenStudy (anonymous):

15.5/100= 155/1000= 31/200 = .155